Grade 11General Physics

calculate the resultant of three forces f1= 50n in a direction 37 degree east of north f2 = 10 n due south and f3 = 50 root 2 in south west direction

sohan
7 Years agoGrade 11

2 Answers

Arun
7 Years ago
Dear student
 
Resultant force in positive y diretion =  50 sin37 + (- 10) – 50sqrt(2) sin45
Resultant force in positive x direction = 50 cos 37 + 0 + – 50 sqrt(2) cos 45
 
solve these As cos 37 and sin 37 are required.
 
Regards
Arun
Khimraj
7 Years ago

Refractive index of slab, μ=1.5

Total apparent length in the glass =0.06+0.04=0.10m

Thickness of the slab, t=μ× Apparent length in the glass =1.5×0.10=0.15m